OpenStudy (anonymous):

The weight of a liquid varies directly with its volume. If the weight of the liquid in a cubical container 4 cm on a side is 128 g, find the weight of the liquid in a cubical container 3 cm on a side.

OpenStudy (anonymous):

Volume, V, of a cube = s^3, where s is the length of one side. V = kW, where k is the constant and W = weight. 4^3 = k(128). k = 4^3 / 128 = 64 / 128 = 1/2 V = (1/2)W. Solve for W for a 5 cm side cubical container. 5^3 = (1/2)W. Multiply both sides by 2. W = 125(2) = 250 g Answer